I have a test coming up. I really need to know if I am on the right track.
1] list the first 4 terms of sequence given by the formula: s[n]=n!/n^2, n>=1.
I came up with 1, 1/2, 2/3, 3/2.
2} list the first 4 terms of sequence given by the recursion formula
a[n]=2a[n-1]+3, with a[1]=1. I came up with (1,3,7,15)
3] the sum of the first 500 terms of aritmetic series 1+4+7+...
I came up with 125,000

Answer by stanbon(75887)

1] list the first 4 terms of sequence given by the formula: s[n]=n!/n^2, n>=1.
I came up with 1, 1/2, 2/3, 3/2.
Correct
---------------------
2} list the first 4 terms of sequence given by the recursion formula
a[n]=2a[n-1]+3, with a[1]=1. I came up with (1,3,7,15)
a(2)=2(1)+3=5
a(3)=2(5)+3=13
a(4)=2(13)+3=29
----------------------
3] the sum of the first 500 terms of aritmetic series 1+4+7+...
I came up with 125,000
a(1)=1
d=3
a(500)=1+499*3=1498
S(500)=(500/2)(1+1498)
S(500)=250(1499)=374,750
------------------------
Cheers,
Stan H.